OpenStudy (anonymous):

-2x + y = 4? how would this look like in a graph?

OpenStudy (anonymous):

Turn that problem into an y = mx + b equation then the graphing is so much easier...

OpenStudy (anonymous):

so it would b y=-2x+4?

jimthompson5910 (jim_thompson5910):

more like y = 2x+4

OpenStudy (anonymous):

but i thought it was negative 2x

OpenStudy (anonymous):

Thanks, Jim ;D

OpenStudy (anonymous):

why is it not negative 2x?

jimthompson5910 (jim_thompson5910):

you add 2x to both sides to isolate y and yw LizzyLove<3

OpenStudy (anonymous):

im lost

jimthompson5910 (jim_thompson5910):

-2x + y = 4 -2x + y + 2x = 4 + 2x ... Add 2x to both sides. (-2x+2x) + y = 4 + 2x 0x + y = 4 + 2x y = 4 + 2x y = 2x + 4

OpenStudy (anonymous):

oh! now i see!!! thank you:)

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

Jim is the smartest person, he knows like EVERYTHING!! He even has hilarious lines... @Juliana16

OpenStudy (anonymous):

yes he has now help me in like 3 problems iv had!

OpenStudy (anonymous):

I think he's too smart. He needs to give some of his smartness to me ;P

jimthompson5910 (jim_thompson5910):

ok i'll try *tries to hit head so some of my 'smartness' leaks out* lol idk

jimthompson5910 (jim_thompson5910):

probably not the smartest way to share brain power...but that's because i hit my head, so don't judge me

OpenStudy (anonymous):

Lol, don't bang your head, dork!! LOl ;D

OpenStudy (anonymous):

lol! funny

jimthompson5910 (jim_thompson5910):

ok i'll spare my head for now

OpenStudy (anonymous):

For now... lol

jimthompson5910 (jim_thompson5910):

only time will tell

jimthompson5910 (jim_thompson5910):

do you know why only time will tell

OpenStudy (anonymous):

Wow, jim. no i dont...

jimthompson5910 (jim_thompson5910):

because time is a snitch

jimthompson5910 (jim_thompson5910):

i guess this is what happens when i get bored lol i go off on random tangents

OpenStudy (anonymous):

Alright, Have a great day/night. Lol jim. @jim_thompson5910 @Juliana16